OpenStudy (anonymous):

X is a factor of 35. Y is a factor of 16. Which number below could not be a value for xy?

OpenStudy (aum):

Factors of 35: 1, 5, 7, 35 (X is one of these numbers) Factors of 16: 1, 2, 4, 8, 16 (Y is one of these numbers) If you provide the choices we can see which ones would fit XY and which ones won't.

OpenStudy (anonymous):

We'll my answers are 10, 24, 40, 56, 70

OpenStudy (dumbcow):

5*2 = 10 5*8 = 40 7*8 = 56 35*2 = 70 can you find 2 numbers that multiply to 24?

OpenStudy (anonymous):

No. I got. It. Thank you :)